A 40 mL solution is 80% acid. How much water
must be added to produce a solution that is 60%
acid?

Answer by stanbon(75887)

A 40 mL solution is 80% acid. How much water
must be added to produce a solution that is 60%
----------------------
Equation:
acid + acid = acid
0.80*40 + 0*x = 0.60(40+x)
Multiply thru by 100 to get:
80*40 + 0 = 60*40 + 60x
20*40 = 60x
x = (1/3)*40
x = 13 1/3 mL (amt. of water needed in the mixture)
===========
Cheers,
stan H.